The legs of an isosceles right triangle each measure 10 inches. What
is the length of the hypotenuse of this triangle, to the nearest tenth of
an inch?

Answer by stanbon(75887)

The legs of an isosceles right triangle each measure 10 inches. What
is the length of the hypotenuse of this triangle, to the nearest tenth of
an inch?
----
hypotenuse = sqrt[10^2 + 10^2] = sqrt(200) = 10sqrt(2)
----
= 10*1.414 = 14.1 inch
======================
cheers,
Stan H